For the system x + y = 2 and y = - x1/4^2 + 3, Could you solve the system? Thanks.
Answer by Alan3354(69443)

x + y = 2
y = - x1/4^2 + 3
------
y = 2 - x
y = - x/16 + 3
----
2 - x = 3 - x/16
32 - 16x = 48 - x
-15x = 16
x = -16/15